OpenStudy (anonymous):

find the simultaneous solution m=n/3 5m/2-5n=20

jimthompson5910 (jim_thompson5910):

The first thing I would do is multiply both sides of 5m/2-5n=20 by 2 This is so you can clear out the fraction

jimthompson5910 (jim_thompson5910):

5m/2-5n=20 2*(5m/2-5n)=2*20 2*(5m/2) + 2*(-5n) = 2*20 5m - 10n = 40 making sense so far?

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

m=n/3 turns into 3m = n when you multiply both sides by 3

jimthompson5910 (jim_thompson5910):

so n = 3m

jimthompson5910 (jim_thompson5910):

5m - 10n = 40 5m - 10(3m) = 40 ... replace n with 3m (works because n = 3m) 5m - 30m = 40 what do you get when you solve for m?

OpenStudy (anonymous):

so u divide 40 by 25

jimthompson5910 (jim_thompson5910):

-25

jimthompson5910 (jim_thompson5910):

because 5-30 = -25

OpenStudy (anonymous):

so then it would be -1.6

jimthompson5910 (jim_thompson5910):

correct, which is the fraction m = -8/5

jimthompson5910 (jim_thompson5910):

once you know the value of m, you can find the value of n

jimthompson5910 (jim_thompson5910):

use n = 3m and replace m with -1.6 or -8/5 and evaluate

OpenStudy (anonymous):

how would it be -8/5

jimthompson5910 (jim_thompson5910):

well using a calculator, you get -8/5 = -1.6 or you can think of -1.6 as -16/10 and reduce that to get -8/5

OpenStudy (anonymous):

oh mmk thank u so much

jimthompson5910 (jim_thompson5910):

-1.6/1 = -16/10 ... multiply top and bottom by 10 to move the decimal over

jimthompson5910 (jim_thompson5910):

np